OpenStudy (anonymous):

What is the radius of the circle whose equation is x^2 + y^2 = 8?

OpenStudy (ash2326):

Standard equation of a circle with center at origin and radius r \[x^2+y^2=r^2\] here we have the equation as \[x^2+y^2=8\] Compare the two equation and find the radius. Can you do that? @Cherishyou

OpenStudy (anonymous):

64

OpenStudy (ash2326):

Look carefully r^2=8, find r from that :)

OpenStudy (anonymous):

dont you sqaure the 8 ?

OpenStudy (ash2326):

We have the equation as \[x^2+y^2=8\] Standard equation with radius r \[x^2+y^2=r^2\] Compare both so \[x^2=x^2\] \[y^2=y^2\] \[r^2=8\]

OpenStudy (anonymous):

2 square root of 2

OpenStudy (ash2326):

yes:D

OpenStudy (anonymous):

Thank Youuu . I Was Confused For A Minute ( : lol

OpenStudy (ash2326):

Np:D